Aluminium displaces hydrogen from dilute HCl whereas silver does not. The E.M.F. of a cell prepared by combining Al/Al+3 and Ag/Ag+ is 2.46 V. The reduction potential of silver electrode is + 0.80 V. The reduction potential of aluminium electrode is: 1.66 V ?1.66 V 3.26 V ?3.26 V

  •  Aluminium displaces hydrogen from dilute HCl whereas silver does not. The E.M.F. of a cell prepared by combining Al/Al+3 and Ag/Ag+ is 2.46 V. The reduction potential of silver electrode is + 0.80 V. The reduction potential of aluminium electrode is:
  • 1.66 V
  • ?1.66 V
  •  3.26 V
  • ?3.26 V

Grade:8

1 Answers

User
6 years ago
-1.66......since aluminium displaces H+ from HCl and age doesn`t.... Hence Al is a reducing agent then,E•cell =E•cathode-E•anode.........................,............................... . where cathode is silver metal ,,, and anode is aluminium, because reducing agent(Al) is always can anode in galvanic cell........... Then 2.46=0.8-xx=0.8-2.46=... -1.66:))